OpenStudy (anonymous):

x^2-2x-5>orequal to 3

OpenStudy (anonymous):

{{x >= 4}, {x <= -2}}

OpenStudy (anonymous):

start with \[x^2-2x-8\geq 0\] then find the zeros. since this is a parabola facing up with will be positive outside them

myininaya (myininaya):

(x-4)(x+2)>=0

OpenStudy (saifoo.khan):

\[x \le -2 \] \[x \ge +4\]

OpenStudy (anonymous):

\[(x-4)(x+2)\geq 0\]wow we all got this one!

OpenStudy (anonymous):

inf?

myininaya (myininaya):

oops i forgot to be brackets

myininaya (myininaya):

(-inf,-2] U [4,inf)

OpenStudy (saifoo.khan):

\[\Huge \infty\]

OpenStudy (anonymous):

\[(-\infty,-2)\cup (4,\infty)\]

myininaya (myininaya):

you mean to include -2 and 4

OpenStudy (anonymous):

\[\Huge \frac{\infty}{\color{blue}{\text{saifoooooo!}}}\]

myininaya (myininaya):

\[(-\infty,-2] \cup [4,\infty)\]

OpenStudy (anonymous):

yeah yeah that is the second time i made that mistake tonight

OpenStudy (anonymous):

\[(-\infty,-2] \cup [4,\infty)\]

myininaya (myininaya):

i win :)

OpenStudy (anonymous):

as per usual

myininaya (myininaya):

lol